OpenStudy (anonymous):

The surface area S of a cube with edge length x is given by S(x) = 6x2 for x > 0. Suppose the cubes your company manufactures are supposed to have a surface area of exactly 42 square centimeters, but the machines you own are old and cannot always make a cube with the precise surface area desired. Write an inequality using absolute value that says the surface area of a given cube is no more than 3 square centimeters away (high or low) from the target of 42 square centimeters.

OpenStudy (anonymous):

\[ \left| 6 x^2 -42 \right |\le 3 \]

OpenStudy (anonymous):

This will give you \[ 42-3= 39 < 6 x^2 < 45=42+3 \]

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

The answer is the first post.

OpenStudy (anonymous):

how did you get there?

OpenStudy (anonymous):

That is what they are asking you to do.

OpenStudy (anonymous):

I see so the absolute value of what S(x)-42(what they want) should be less than or equal to 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i can see clearly now. Thank you so much. Nobody wanted to answer it.

OpenStudy (anonymous):

yw